Hellllllpppppppp
On Monday, Archie paid $2.54 for three doughnuts and two coffees. On Tuesday he paid $2.46 for two doughnuts and three coffees. On Wednesday he was tired of paying the tab and went out for coffeeby himself. What was his bil for one doughnut and one coffee

Answer by Paul(988)

Let the amount of coffe be x
Let the amount of donuts be y
Now 3 donuts and 2 coffe cost 2.54
Expression: 3y+2x=2.54
Solve for x
2x=2.54-3y
x=1.27-1.5y (subsitution)

Now 2 donuts (y) and 3 coffee (x) cost 2.46
2y+3x=2.46
Subsitute
2y+3(1.27-1.5y)=2.46
2y-4.5y+3.81=2.46
-2.5y=-1.35
y=0.54

x=1.25-1.5(0.54)
x=0.44

Hence, a coffee cost about $0.54, and a donut costs $0.44
Paul.